OpenStudy (mattduhhelper):

Which expressions are equivalent to −4.5+(−5.5)+(−3.1)−4.5+(−5.5)+(−3.1) ? Select each correct answer. −(4.5+5.5)+(−3.1)−(4.5+5.5)+(−3.1) −4.5−(5.5+3.1)−4.5−(5.5+3.1) −4.5−(5.5−3.1)−4.5−(5.5−3.1) −(4.5+5.5+3.1)

OpenStudy (alivejeremy):

−4.5+(−5.5)+(−3.1)−4.5+(−5.5)+(−3.1)=−26.2

OpenStudy (alivejeremy):

−(4.5+5.5)+(−3.1)−(4.5+5.5)+(−3.1)=−26.2

OpenStudy (alivejeremy):

−4.5−(5.5+3.1)−4.5−(5.5+3.1)=−26.2

OpenStudy (alivejeremy):

−4.5−(5.5−3.1)−4.5−(5.5−3.1)=−13.8

OpenStudy (mattduhhelper):

Thank you goku fan! :D <3

OpenStudy (alivejeremy):

−(4.5+5.5+3.1)=−13.1

OpenStudy (alivejeremy):

You're welcome! I'm a fan of Vegito